Lupe made a down payment of $2000 toward the purchase of a new car. To pay the balance of the purchase price, she has secured a loan from her bank at the rate of 10%/year compounded monthly. Under the terms of her finance agreement she is required to make payments of $230/month for 48 months. What was the cash price of the car? (Round your answer to the nearest cent.)
Answer #1

The present value of loan amount excluding down payment amount is:

=PV(rate,nper,pmt)

=PV(10%/12,48,-230)

=9068.48

Total cash price =9068.48+2000 =11,068.48